Jacob Lee invested $600 in a savings account paying 8% interest compounded twice a year. What will be his interest at the end of the year? Round to the nearest dollar.

$48

$24

$49- Not the answer

$25
Answer #1

Interest for first 6 month = 600*4% = $24

Interest for next 6 month = (600+24)*4% = $24.96

Total interest for the year = 24+24.96 = 48.96 or rounded to $49

If 49 is not thd correct answer the try $25.